OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (mathstudent55):

Multiply the first equation by 7 Multiply the second equation by 3 Then subtract the new second equation from the new first equation. Then solve for y.

OpenStudy (anonymous):

wait i got x-27=54 +27 81 x=

OpenStudy (mathstudent55):

No. Here's the first equation: 3x - 6y = 3 Can you multiply the entire equation by 7? That means multiply every term by 7.

OpenStudy (anonymous):

ya 21x-42=21

OpenStudy (anonymous):

then what

OpenStudy (mathstudent55):

I wrote all the instructions above. Just follow them.

OpenStudy (mathstudent55):

Look at my first response.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Don't forget the y with the 42: 21x-42y=21

OpenStudy (anonymous):

thats what i got for the first one but i got this for yhe second one:21x-15=-33

OpenStudy (mathstudent55):

No it's not. You left the y out of the first one, and now you left the y out of the second one.

OpenStudy (rmrjr22):

you want to cancel out either a y value or an x value...

OpenStudy (mathstudent55):

It should read: 21x - 15y = -33

OpenStudy (anonymous):

ya igot that one

OpenStudy (mathstudent55):

First you just multiply the entrie equations by the numbers. Leaving out the variable y does not mean you canceled it. It means you have an error. You can't just leave y out.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

here are the two correct equations after multiplying: 21x - 42y = 21 21x - 15y = -33 Now subtract the second equation from the first one. 21x - 21x - 42y - (-15y) = 21 - (-33)

OpenStudy (anonymous):

ok see it

OpenStudy (anonymous):

i got x-27=21-(-33)

OpenStudy (mathstudent55):

What is 21x - 21x?

OpenStudy (mathstudent55):

21x - 21x - 42y - (-15y) = 21 - (-33) 0 -27y = 54

OpenStudy (mathstudent55):

Now divide both sides by -27 to get y

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

thankyou

OpenStudy (mathstudent55):

wlcm